OpenStudy (anonymous):

Select the expression below that is identical to \[\frac{ 1 }{ 1-\cos x }+\frac{ 1 }{ 1+\cos x }\] A.)\[2\sec ^{2}\theta\] B.)1 C.)\[2\cos ^{2} x\] D.)\[2\sin ^{2} x\] E.)\[2\csc ^{2} x\]

OpenStudy (whpalmer4):

Do you know how to make a common denominator so you can add those two fractions together? After you do so, I think you'll see something you recognize from your trig identities...

OpenStudy (anonymous):

yes. But i'm not sure I can handle the mess that ensues.

OpenStudy (whpalmer4):

Sure you can...

OpenStudy (anonymous):

ALL RIGHT. Lets give it a try! :)

OpenStudy (whpalmer4):

Here, let me suggest a substitution that will make it less scary-looking :-) Let \(a = \cos x\) so we can rewrite it as \[\frac{1}{1-a}+\frac{1}{1+a}\]Now we multiply the left fraction top and bottom by the denominator of the right fraction, and the right fraction top and bottom by the denominator of the left fraction. Do that and tell me what you get

OpenStudy (anonymous):

\[\frac{ 1(1+\cos x) }{ (1-\cos x)(1+\cos x) }+\frac{ 1(1- \cos x) }{ (1-\cos x)(1+\cos x) }\]

OpenStudy (anonymous):

I already started before you suggested the "a" idea.

OpenStudy (anonymous):

So now you just add then simplify.

OpenStudy (whpalmer4):

Right. Either before or after you add, I'd like you to multiply out the denominator...

OpenStudy (anonymous):

what do you mean?

OpenStudy (whpalmer4):

You've got (1-a)(1+a) — what is that after you expand it?

OpenStudy (whpalmer4):

(1-cos x)(1+cos x) if you prefer

OpenStudy (anonymous):

1 + cos x ^2

OpenStudy (whpalmer4):

close, isn't it 1 - cos^2 x?

OpenStudy (anonymous):

yes. Shame upon my family. :)

OpenStudy (anonymous):

So thats the both the bottom and the top portions so it equals one.

OpenStudy (whpalmer4):

it happens. easier to see what you're doing with my trick, which is why I like it. sin^x + cos^x = 1 1 - cos^x = sin^x (that's the trig identity I was hoping you would remember) so what does the numerator simplify to?

OpenStudy (anonymous):

okay okay so the denominator = sin^2, the numerator would be 1(1+a)+1(1-a) which is just (1+a)+(1-a)

OpenStudy (whpalmer4):

and that simplifies to...

OpenStudy (anonymous):

1

OpenStudy (anonymous):

If i'm not retarded which is possible so...

OpenStudy (whpalmer4):

1 + a + 1 - a = ?

OpenStudy (whpalmer4):

we could rearrange the order to 1 + 1 + a - a = 1 + 1 = 2

OpenStudy (anonymous):

Here i was overthinking the whole thing. Thats embarrassing.

OpenStudy (anonymous):

okay so its 2/ sin ^2

OpenStudy (whpalmer4):

so, that gives us \[\frac{2}{sin^2 x}\] but that isn't on our list of choices, is it?

OpenStudy (anonymous):

nope.

OpenStudy (whpalmer4):

they throw in 2 sin^x on the list as a trap for the unwary, but we're too smart for that! do you remember the definition of secant and cosecant? (sec x, csc x)

OpenStudy (whpalmer4):

one is the reciprocal of sin x, the other is the reciprocal of cos x. boy, wouldn't it be handy if they had matching first letters to help you remember which is which :-) they don't, however.

OpenStudy (whpalmer4):

\[\sec x = \frac{1}{\cos x}\]\[\csc x = \frac{1}{\sin x}\]

OpenStudy (anonymous):

so we have 2 csc ^2 x which is answer E

OpenStudy (anonymous):

I just need like a super list of all the identities.

OpenStudy (anonymous):

You were awesome dude, thanks for helping me out. I actually understood it :)

OpenStudy (whpalmer4):

you got it! yeah, a lot of it is just practice, and learning to recognize the identities in different forms...

OpenStudy (whpalmer4):

I'll confess I had to look up whether sec or csc was the one we wanted :-)

OpenStudy (whpalmer4):

anyhow, remember my substitution trick. it works well with expressions that have lots of square roots and so on in them, too. if nothing else, you save some time writing :-)

OpenStudy (anonymous):

Okay sounds like a deal. :) I'll try to figure out the next one on my own before posting it so expect me back soon! :)

OpenStudy (whpalmer4):

for extra speed, remember that whenever you have fractions like you have in this one where the denominators are x-y and x +y that those are the factors of x^2 -y^2 and you can just right there

OpenStudy (anonymous):

thats pretty sweet. thanks! You're just stuffed with math tips. I'm more of a writer so I don't get this math stuff. :)